OpenStudy (kjay):

If y varies inversely as x, and y =3 when x = 6, find the value of x when y = 18

OpenStudy (anonymous):

aaahhh i hate these kind of questions :P

OpenStudy (kjay):

same here. :b

OpenStudy (anonymous):

k the formula for "y varies inversly with x" is \[y = k / x\]

OpenStudy (anonymous):

so first you put in the values they told you about (y=3 and x=6) and solve for k

OpenStudy (anonymous):

once you know k you set up another equation (using k=? and y =18)

OpenStudy (anonymous):

am i making sense? :)

OpenStudy (kjay):

yes(:

OpenStudy (anonymous):

so you get it?? :D

OpenStudy (kjay):

yes I do, thanks much! (:

OpenStudy (anonymous):

k work it out and tell me what you get :)

OpenStudy (kjay):

hmm. I got 18. Which I think is incorrect. :/

OpenStudy (lgbasallote):

for someone who hates these kinds of questions you're doing well @yummydum :p haha

OpenStudy (anonymous):

nope k=18 is right :)

OpenStudy (anonymous):

thanks @lgbasallote :D

OpenStudy (kjay):

really? oh! awesome! thanks! (:

OpenStudy (anonymous):

so now you do \[18 = 18/x\]

OpenStudy (anonymous):

and get......?

OpenStudy (kjay):

2! :D

OpenStudy (anonymous):

uh wat?

OpenStudy (kjay):

18 = 2

OpenStudy (anonymous):

try that again :)

OpenStudy (anonymous):

solve for x : 18=18/x

OpenStudy (kjay):

18/x = 18/6 = 3

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

the only info you have is 18=18/x...now solve for x

OpenStudy (kjay):

ohhhh

OpenStudy (anonymous):

so you would do 18=18/x......and get 18/18=x ....so x=? :D

OpenStudy (kjay):

1?

OpenStudy (anonymous):

YAYY! :D

OpenStudy (kjay):

thanks! :D

OpenStudy (anonymous):

no problem kjay :) ....msg me if you have any other questions!! :)

OpenStudy (kjay):

will do! (:

OpenStudy (anonymous):

p.s....fanning me might just help :P

OpenStudy (anonymous):

k cool :D

OpenStudy (kjay):

I was just about to do that (x

OpenStudy (anonymous):

your amazing! :D

OpenStudy (kjay):

no, you're amazing! :D

OpenStudy (anonymous):

oh shucks! :P